These bags have a mean weight of 10kg with a standard deviation of 2kg.

a) Give the point estimate for the mean weight of all bags of cement in the warehouse. [ 1 marks ]

b) Compute the standard error. [1 marks ]

c) Construct the confidence interval of the mean weight of all the bags of cement in the warehouse

using 90% confidence level. [ 4 marks ]

d) If the inspector needs to estimate the mean weight of the bags with a maximum error of 0.5kg at the 90% confidence level, how large a sample will he need to take


1
Expert's answer
2021-06-08T17:32:45-0400

a) The sample mean "\bar{x}" is a point estimate of the population mean, "\mu."  

 The point estimate for the mean weight of all bags of cement in the warehouse is 10 kg.


b) "SE_{\bar{x}}=s/\sqrt{n}"

Suppose we take 40 bags of cement in the warehous. Then the standard error is


"SE_{\bar{x}}=\dfrac{s}{\sqrt{n}}=\dfrac{2}{\sqrt{40}}\approx0.3162"

c) Suppose we take 40 bags of cement in the warehous.

The critical value for "\alpha=0.1" and "df=n-1=40-1=39" degrees of freedom is "t_c=1.684875."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(10-1.684875\times\dfrac{2}{\sqrt{40}},10+1.684875\times\dfrac{2}{\sqrt{40}})"

"\approx(9.4672, 10.5328)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "9.4672<\mu<10.5328," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(9.4672, 10.5328)."


d) "\alpha=0.1, s=2"


"z_{1-\alpha/2;n-1}\times\dfrac{s}{\sqrt{n}}\leq0.5"

"n\geq(2z_{1-\alpha/2;n-1}\times s)^2"

"n\geq(4z_{1-\alpha/2;n-1})^2"

"n=45"


"45\geq(4\cdot1.68023)^2, False"

"n=46"


"46\geq(4\cdot1.679427)^2, True"

He will need to take 46 bags.